Use the information in the following table to determine the percentage changes in the U.S. and French real minimum wages between 1957 and 2016.

United States

France

Year

Minimum wage

​ (dollars per​ hour)

CPI

Minimum wage

​(euros per​ hour)

CPI

1957

​$1.00

27

0.19 euros

8

2016

​$7.25

240

9.76 euros

106

​Sources: For​ 1957: John M.​ Abowd, Francis​ Kramarz, Thomas​ Lemieux, and David N.​ Margolis, "Minimum Wages and Youth Employment in France and the United​ States," in D. Blanchflower and R.​ Freeman, eds., Youth Employment and Joblessness in Advanced

Countries​,

​Chicago: University of Chicago​ Press, 1999, pp.​ 427-472 (the value for the minimum wage is given in​ francs; it was converted to euros at a conversion rate of 1 euro​ = 6.55957​ francs); For​ 2016: INSEE online data​ bank, www.insee.fr; U.S. Department of​ Labor; and U.S. Bureau of Labor Statistics.

Calculate the percentage change in U.S. real minimum wages between 1957 and​ 2016:

-----------

​(Enter your response as a percentage rounded to two decimal​ places).
Answer #1

Real minimum wages in US in 1957 = (1/27) x 100 = $3.704

Real minimum wages in US in 2016 = (7.25/240) x 100 = $3.021

% change in real wages = [(3.021-3.704)/3.704] x 100 = -18.44%

Thus, real minimum wages in US fell by 18.44% between 1957 and 2016

Note:

Formula used:

Real w = (Nominal w / CPI) x 100

% change = [(New value - Old value) / Old value ] x 100